Solve the system of equations by using substitution. Express the answer as an ordered pair:
x+5y=2
x-1=2y

Answer by Boreal(15235)

x+5y=2
x-1=2y
x=2y+1
Substitute into the first
2y+1+5y=2
7y=1
y=(1/7)
x=2(1/7)+1=(9/7)
try this in the first
9/7=(2/7)+1 true
( (9/7), (1/7) )